# Learning Distances from Data with Normalizing Flows and Score Matching

### v1

@@ A.2 Relaxation scheme
By substituting this into the differential equation we have
@@ Appendix B Experimental Details
Figure 5: The 5 two-dimensional datasets.
Software libraries
We used the numpy [Harris et al., 2020] and scipy [Virtanen et al., 2020] Python libraries for basic numerical and graph operations and matplotlib [Hunter, 2007] for plotting. We used scikit-learn [Pedregosa et al., 2011] to compute nearest neighbor graphs and pytorch [Paszke et al., 2019] for training neural networks. We used FrEIA [Ardizzone et al., 2018-2022] for constructing normalizing flows.
Two-dimensional datasets
@@
  • Standard normal
  • Gaussian mixture with 3 components
  • Circle
  • 1 spiral
  • 2 spirals
Normalizing flow training
We used the FrEIA library [Ardizzone et al., 2018-2022] to design normalizing flows using spline coupling blocks [Durkan et al., 2019]. We used the following hyperparameters:
@@
Score model training
We use fully connected networks to predict the score, with 6 hiddens layer with the width depending on the dimension of the input. We use the following hyperparameters:
  • Hidden layer width (as a function of input dimension $D$)
@@ Appendix C Additional experimental results
Figure 6: Increasing the number of nearest neighbors does not significantly improve the convergence of the power-weighted graph method on two-dimensional datasets.
C.1 Convergence of alternative graph edge weightings
Figure 7: Alternative edge weightings based on the nearest neighbor density estimator do not converge. The same weightings based on the ground truth density do converge. Experiments performend on the GMM two-dimensional dataset.
@@ C.1 Convergence of alternative graph edge weightings
We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge. See Fig. 7.
Figure 8: Detailed plot of convergence behavior in higher dimensions, unscaled metric. We see that all graph-based methods plateau in performance, whereas the relaxation-based method continues to show good convergence.
Figure 9: Detailed plot of convergence behavior in higher dimensions, scaled metric. We see that all graph-based methods plateau in performance, whereas the relaxation-based method continues to show good convergence.

### v2

@@ A.2 Relaxation Scheme
By substituting this into the differential equation (Equation 21) we have
@@
A.3 Justification of Scaling Fermat Distances with Dimension
We propose to scale the metric using a temperature equal to the dimension of the data. In order to justify this, consider the behavior of a generalized Gaussian distribution of the form $p(x)\propto\exp(-\|x\|^{\alpha}/\alpha)$. It can be shown that $\mathbb{E}[\|x\|^{\alpha}]=D$ and thus the typical distance from the origin of points sampled from $p$ scales as $D^{1/\alpha}$. The score in this case is $s(x)=-\|x\|^{\alpha-2}x$.
Due to the rotational symmetry of the distribution, we can suppose without loss of generality that a sample $x$ lies in the plane of the first two coordinates and hence the problem reduces to the 2-dimensional one. However, since the typical distance of sampled points scales with the dimension, in order to avoid very exaggerated geodesics in high dimensions which deviate sharply towards the origin, we need to scale the density appropriately. By rescaling $\gamma$ in Equation 3 by $D^{1/\alpha}$, the score scales like $D^{(\alpha-1)/\alpha}$, and so the terms in the equation scale as:
$\ddot{\gamma}-2\beta(s\cdot\dot{\gamma})\dot{\gamma}+\beta s\|\dot{\gamma}\|^{2}$ $\to D^{1/\alpha}\ddot{\gamma}-2\beta D^{(\alpha+1)/\alpha}(s\cdot\dot{\gamma})\dot{\gamma}+\beta D^{(\alpha+1)/\alpha}s\|\dot{\gamma}\|^{2}$ (32)
$=D^{1/\alpha}\bigl{(}\ddot{\gamma}-2\beta D(s\cdot\dot{\gamma})\dot{\gamma}+\beta Ds\|\dot{\gamma}\|^{2}\bigr{)}$ (33)
Therefore the geodesic equation can scale consistently if, and only if, $\beta=1/D$. The same reasoning applies to the reparameterized Equation 4.
We can apply a similar argument to the Student-t distribution of the form $p(x)\propto(1+\|x\|^{2}/\nu)^{-(\nu+D)/2}$ with $\nu>2$. In the limit $\nu\to\infty$, the distribution becomes normal, and otherwise is characterized by heavier tails. The expected squared norm $\mathbb{E}[\|x\|^{2}]=\nu D/(\nu-2)$, so the typical distance from the origin scales like $\sqrt{D}$, as in a Gaussian distribution. The score function has the form $-\tfrac{\nu+D}{\nu+\|x\|^{2}}x$, and with $x\sim\sqrt{D}$, we have $s\sim x\sim\sqrt{D}$, again the same behavior as the Gaussian. Hence the scaling behavior is approximately the same as the Gaussian distribution, so the above argument for $\beta=1/D$ applies as well. Given that the same scaling behavior holds as in the generalized Gaussian, it suggests that $\beta=1/D$ is a natural default choice across a wide range of distributions.
@@ Appendix B Experimental Details
Figure 8: Samples from the 5 two-dimensional datasets.
B.1 Software Libraries
We used the numpy (Harris et al., 2020) and scipy (Virtanen et al., 2020) Python libraries for basic numerical and graph operations and matplotlib (Hunter, 2007) for plotting. We used scikit-learn (Pedregosa et al., 2011) to compute nearest neighbor graphs and pytorch (Paszke et al., 2019) for training neural networks. We used FrEIA (Ardizzone et al., 2018-2022) for constructing normalizing flows.
B.2 Two-Dimensional Datasets
@@
  1. 1.
    Standard normal distribution
  2. 2.
    Gaussian mixture with 3 components:
    • Means: $\mu_{1}=(2,1.4)$, $\mu_{2}=(6.5,6.3)$, $\mu_{3}=(8,1)$
    • Covariances: $\Sigma_{1}=\begin{pmatrix}3&2.5\\ 2.5&3\end{pmatrix}$, $\Sigma_{2}=\begin{pmatrix}3&0\\ 0&3\end{pmatrix}$, $\Sigma_{3}=\begin{pmatrix}2&-0.8\\ -0.8&2\end{pmatrix}$
    • Weights: $(0.25,0.5,0.25)$
  3. 3.
    Circle with radius 1 and Gaussian noise ($\sigma=0.08$)
  4. 4.
    Single spiral with 1.75 rotations ($3.5\pi$ radians) and Gaussian noise ($\sigma=0.05$)
  5. 5.
    Two interleaved spirals with 1 rotation each and Gaussian noise ($\sigma=0.045$)
The spirals reach a radius of 1 after a single rotation.
B.3 Normalizing Flow Training
We used the FrEIA library (Ardizzone et al., 2018-2022) to design normalizing flows using spline coupling blocks (Durkan et al., 2019). We used the following hyperparameters:
@@ B.3 Normalizing Flow Training
B.4 Score Model Training
We use fully connected networks to predict the score, with 6 hidden layers of varying width depending on the input dimension. The architecture consists of:
  • Input layer → 6 hidden layers → Output layer (dimension matching input)
  • Hidden layer widths:
@@ Appendix C Additional Experimental Results
C.1 MNIST experiments
Table 1: For each digit, we compute the mean and standard deviation of the log distances between the cluster mean and 20 randomly sampled points of that class (intra-class log distance), as well as between the cluster mean and 20 randomly sampled points of other classes (inter-class log distance). We also compute the difference, which gives an indication of how well-separated the digits are in terms of log distance.
Digit Intra-Class Log Distance Inter-Class Log Distance Difference (Inter–Intra)
0 2.41 ± 0.40 3.73 ± 0.27 1.32 ± 0.49
1 1.69 ± 0.49 3.73 ± 0.31 2.04 ± 0.58
2 3.56 ± 0.29 4.05 ± 0.19 0.49 ± 0.35
3 3.03 ± 0.33 3.78 ± 0.24 0.75 ± 0.41
4 3.04 ± 0.38 3.77 ± 0.36 0.73 ± 0.53
5 3.43 ± 0.35 4.02 ± 0.18 0.59 ± 0.39
6 2.89 ± 0.41 3.79 ± 0.31 0.91 ± 0.51
7 2.86 ± 0.47 3.81 ± 0.30 0.95 ± 0.56
8 3.29 ± 0.33 3.74 ± 0.21 0.46 ± 0.39
9 2.76 ± 0.45 3.56 ± 0.38 0.80 ± 0.59
Figure 9: Each row visualizes an MNIST geodesic, obtained by relaxation (Algorithm 1) using a learned score model within a 10-dimensional autoencoder latent space.
We train an autoencoder to compress MNIST (LeCun et al., 2002) digits to a 10-dimensional space and subsequently train a sliced score matching model (Song et al., 2020) and normalizing flow (Dinh et al., 2017) on the latent distribution. Both encoder and decoder use fully connected layers with ReLU activations, with a single hidden layer of width 400. The score model has input and output dimension 10 and has hidden layers of width 128, 256, and 128, with SiLU activations. The normalizing flow is built with the FrEIA framework (Ardizzone et al., 2018-2022), made up of 4 affine coupling blocks, with a permutation between each block. Each block uses a coupling network with 2 hidden layers of width 128 and SiLU activations. All models are trained for 400 epochs with the Adam optimizer, the autoencoder with learning rate $10^{-3}$ and the score model and normalizing flow with learning rate $10^{-4}$.
To compute distances between two points sampled from the latent distribution, we the same procedure as in the main text: i) construct a density-weighted graph (Algorithm 2) from the normalizing flow model ii) use Dijkstra’s algorithm to find the shortest path between the points according to the graph iii) solve the relaxation problem (Algorithm 1) with the score model, using the graph-based path as initialization iv) evaluate the length of the relaxed path using the normalizing flow model, according to Equation 2.
In Table 1 we report mean and standard deviation of log distances between each digit cluster mean and 20 random samples of the same or other digits. We see that same-class distances are always lower than other-class distances, up to a safe margin of error. We also make the following observations: 1s cluster tightly, likely due to having fewer active pixels and therefore having higher likelihoods; 8s are more ambiguous, likely due to easy deformations into other digits (it is quite easy to remove some pixels to turn an 8 into a 3, 5, 6 or 9). We leave deeper analysis of Fermat distances on MNIST to future work.
In Figure 9 we visualize geodesics between 5 randomly sampled pairs from the latent distribution. The geodesics are solved in the latent space and the resulting latent points are decoded to image space for visualization. We see that the geodesics follow high-density regions of the distribution, leading to interpretable and reasonable transitions between the digits.
C.2 Convergence of Alternative Graph Edge Weightings
Figure 10: Alternative edge weightings based on the nearest neighbor density estimator converge poorly. The same weightings based on the ground truth density converge well. Experiments performed on the GMM two-dimensional dataset.
@@ C.2 Convergence of Alternative Graph Edge Weightings
We find that all 4 have very similar performance to the power-weighted graph, whereas the equivalent estimators using the ground truth density quickly converge. See Figure 10.
C.3 Trade-off Between Log Density and Score Estimation
Figure 11: Mean integrated squared error (MISE) for log density and score estimation using KDE with varying bandwidth. The optimal bandwidth for log density estimation is smaller than the optimal bandwidth for score estimation.
We speculate that there is a fundamental trade-off between accurately estimating log densities and their gradients (score functions) that applies to all density estimation methods, including the normalizing flows used in our main experiments. To illustrate this principle in a simple setting where we can compute exact errors, we analyze kernel density estimation (KDE) of the standard normal distribution in one dimension using Gaussian kernels, with 1000 samples drawn from the distribution. While this example is much simpler than the high-dimensional problems and neural network models considered in the main text, it clearly demonstrates how optimizing for density estimation can lead to poor gradient estimates and vice versa.
For a set of samples $\{x_{i}\}_{i=1}^{n}$, the KDE estimate with bandwidth $\sigma$ is:
$\hat{p}(x)=\frac{1}{n\sigma\sqrt{2\pi}}\sum_{i=1}^{n}\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)$ (34)
The log density and score estimates are then:
$\log\hat{p}(x)$ $=-\log(n\sigma\sqrt{2\pi})+\log\sum_{i=1}^{n}\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)$ (35)
$\hat{s}(x)$ $=-\frac{\sum_{i=1}^{n}(x-x_{i})\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)}{\sigma^{2}\sum_{i=1}^{n}\exp\left(-\frac{(x-x_{i})^{2}}{2\sigma^{2}}\right)}$ (36)
Figure 12: Comparison of KDE estimates using different bandwidths. Top row: Log density estimates. Bottom row: Score estimates. Left column: Using the bandwidth that minimizes log density MISE. Right column: Using the bandwidth that minimizes score MISE. Ground truth shown in dotted orange, KDE estimate in solid blue. The optimal bandwidth for log density leads to noisy score estimates, while the optimal bandwidth for score estimation leads to oversmoothed log densities.
As shown in Figure 11, the optimal bandwidth for minimizing the mean integrated squared error (MISE) of the log density estimate is smaller than the optimal bandwidth for minimizing the MISE of the score estimate. This reflects a fundamental trade-off: accurate density estimation requires capturing fine details of the distribution, while accurate score estimation requires smoothing out noise to obtain reliable gradients.
Figure 12 visualizes this trade-off. When using the bandwidth optimal for log density estimation, the score estimates are quite noisy. Conversely, when using the bandwidth optimal for score estimation, the log density estimates are oversmoothed. This observation helps explain why we obtain better results using dedicated score models rather than differentiating density models.